# PAID: A Framework of Product-Centric Advertising Image Design

### v1

@@ Abstract.
In E-commerce platforms, a full advertising image is composed of a background image and marketing taglines. Automatic ad image design reduces human costs and plays a crucial role. For the convenience of users, a novel automatic framework named Product-Centric Advertising Image Design (PAID) is proposed in this work. PAID takes the product foreground image, required taglines, and target size as input and creates an ad image automatically. PAID consists of four sequential stages: prompt generation, layout generation, background image generation, and graphics rendering. Different expert models are trained to conduct these sub-tasks. A visual language model (VLM) based prompt generation model is leveraged to produce a product-matching background prompt. The layout generation model jointly predicts text and image layout according to the background prompt, product, and taglines to achieve the best harmony. An SDXL-based layout-controlled inpainting model is trained to generate an aesthetic background image. Previous ad image design methods take a background image as input and then predict the layout of taglines, which limits the spatial layout due to fixed image content. Innovatively, our PAID adjusts the stages to produce an unrestricted layout. To complete the PAID framework, we created two high-quality datasets, PITA and PIL. Extensive experimental results show that PAID creates more visually pleasing advertising images than previous methods.
Figure 1. Generated advertising images by our methods with only a product subject image, a target size, and taglines as input.
@@ 1. Introduction
Advertising images for products play a crucial role in commercial recommendation systems. Visually harmonious, aesthetically pleasing, and attractive advertising images tend to achieve relatively high click-through rates (Wang et al., 2021). With the rapid advancement of generative methods (Goodfellow et al., 2014; Menick and Kalchbrenner, 2019; van den Oord et al., 2016; Ho et al., 2020; Rombach et al., 2022a), it has become possible to automatically generate advertising images by providing a product foreground image, marketing taglines, and target size as input.
To accomplish this task, by referring to existing approaches (Du et al., 2024; Weng et al., 2024; Wang et al., 2024b; Lin et al., 2023; Zhou et al., 2022), we can first utilize heuristics and predefined rules to generate background text prompts. These prompts are then fed into text-to-image inpainting models to generate backgrounds for product foregrounds, resulting in a composite advertisement image. If marketing taglines are required, the layout of these taglines is predicted based on the generated image, and the taglines are rendered onto the image using techniques like attribute prediction. However, this approach suffers from several limitations: a) it cannot adaptively generate background prompts based on the specific content and shape of the product foreground, leading to poor foreground-background harmony and limited diversity; b) the position of the product foreground is fixed and cannot be adjusted according to the descriptions of the product foreground and background prompt, potentially resulting in inappropriate product size and poor composition; and c) when taglines need to be added, the image content is already finalized, restricting the spatial layout available for tagline placement.
@@
In this paper, we introduce a novel framework called Product-Centric Advertising Image Design (PAID). This framework centers around the integration of product foreground images and marketing taglines to automatically create more harmonious and aesthetic advertising images. Although PAID remains a multistage approach for generating advertising images, we have adjusted the stages and task definitions to optimize existing methods concerning product-background compatibility and the spatial arrangement of product foregrounds and marketing taglines.
As illustrated in Figure 2, we first introduce a task to generate prompts, using a Visual Language Model (VLM) (Dong et al., 2024) to adaptively create foreground and background prompts that better match the scene, placement, and angle based on product foregrounds. Following this, we propose a layout generation task that organizes the arrangement of the graphic elements (marketing taglines and decorative underlays), product foreground, and other nongraphic elements according to the product foreground, background prompt, marketing taglines, and target size. By advancing the layout stage and arranging the product foreground and marketing information concurrently, as opposed to generating the overall image first and then deciding the marketing text layout, the latter becomes less constrained by the image. This allows for a more rational presentation that is suitable for the needs of the content. In this process, the size and position of the product and other nongraphic elements are considered together, which addresses the limitations of adapting the foreground based on prompts in the inpainting task and reduces conflicts between nongraphic elements and marketing information that could adversely affect readability and attractiveness. Next, we train an inpainting text-to-image model based on SDXL (Podell et al., 2024) that can be controlled by input layouts. Combining inpainting, layout, and prompt control effectively is challenging due to potential conflicts among multiple control signals. A LoRA (Hu et al., 2022) adaption strategy is adopted to balance conflicts and facilitate training. Since layout mainly influences higher-level abstract information, we apply layout control only in the deeper blocks of the UNet, combined with a multi-scale training strategy to enhance training stability and efficiency. Finally, we utilize an attribute prediction and graphics rendering module to overlay graphic elements onto the image. To build the framework and verify its effectiveness, we have established a corresponding dataset, collecting a total of 38,017 images with detailed image captions, product foreground images, and layouts of graphic and nongraphic elements.
@@
  • We propose a novel multi-stage framework for advertising image design, which centers on product foregrounds and taglines and enhances the overall content prominence while maintaining the harmony and aesthetics of the images.
@@
  • To arrange product foregrounds and taglines more effectively, a VLM-based layout generator is proposed. The generator can mix and arrange graphic and nongraphic elements with the target size and prompt. Accordingly, an image generation model that simultaneously accepts layout, prompt, and foreground control is proposed. Several strategies are introduced to enhance effectiveness and efficiency.
  • We created two datasets: one with 38,017 annotated images for advertising image design and another with 12,247 annotated images for background generation. We use these datasets to train and validate the models in the framework.
@@ 2.1. Advertising Image Design
In recent years, various algorithms are proposed to automate advertising image design in e-commerce scenario (Zhou et al., 2022; Hsu et al., 2023; Li et al., 2023b; Seol et al., 2024; Lin et al., 2023; Li et al., 2023a). Typically, these approaches can be divided into two subcategories: background image-based and foreground-based methods. Background image-based methods adopt a full image as input and design models to predict the layout, followed by attribute prediction and graphics rendering techniques to construct the final ad image. Here, a layout means the class and position of texts, underlays, or other graphic elements. CGL-GAN (Zhou et al., 2022) and DS-GAN (Hsu et al., 2023) design transformer architectures to leverage the content of the background image and its saliency map for generating layout. RADM (Li et al., 2023b) builds a diffusion model and uses the image content and tagline content to derive the layout. It is the first work to consider tagline content. PosterLlama (Seol et al., 2024) leverages the rich design knowledge in Large Language Models (LLMs) to predict the layout. AutoPoster (Lin et al., 2023) introduces a complete pipeline which first generates a layout based on a background image, and then predicts the attributes of graphic elements and renders them onto the image. For these methods, the spatial layout is restricted due to fixed image content, which further restricts the tagline content, such as number and length.
In contrast, P&R (Li et al., 2023a) is a foreground-based method that leverages both the product foreground image and tagline content to predict the layout of the product foreground and taglines. It subsequently designs a module network to outpaint the background and render the texts according to the given background prompt and generated layout. The tagline content is not limited, and the layout is more flexible. However, P&R ignores the harmony between the background prompt and the generated layout. The layout may conflict with background generation. In our PAID framework, the overall layout of graphic elements, product foreground, and other nongraphic elements is generated jointly according to the product foreground, background prompt, and tagline contents, aiming to achieve the best harmony among them.
@@ 2.2. Controllable Image Generation
Diffusion models (Ho et al., 2020; Rombach et al., 2022a) have significantly improved image generation quality. To control instances and improve image quality for text-to-image models, some methods introduce layout guidance. These layout-to-image methods can be grouped into training-free and training-based categories. Mostly, training-free layout-to-image methods either manipulate the attention map of instances in prompt (Kim et al., 2023), or compute attention-based loss and back-propagate the gradients in de-noising procedure (Chefer et al., 2023; Xie et al., 2023; Chen et al., 2024). MultiDiffusion (Bar-Tal et al., 2023) conducts inference for each instance and then combines their latent vectors. Training-based methods modify the network architectures and fine-tune the diffusion model with layout-image data (Li et al., 2023c; Wang et al., 2024a; Zhou et al., 2024). GLIGEN (Li et al., 2023c) and InstanceDiffusion (Wang et al., 2024a) add a gated self-attention layer to introduce layout information. MIGC (Zhou et al., 2024) conducts per-instance shading and aggregates all the shaded instances.
For advertising image generation (Du et al., 2024; Wang et al., 2023; Cao et al., 2024; Li et al., 2023a), text-to-image models equipped with inpainting capabilities are frequently utilized. The inpainting technique is employed to preserve product characteristics, while the text prompt is used to describe the background. The foreground image is utilized by incorporating extra channels into the UNet architecture (Rombach et al., 2022a) or employing an inpainting ControlNet (Zhang et al., 2023). To better control the overall layout of advertising images and improve the quality of generated images, we combine inpainting with layout control. Currently, there is limited exploration in this area. SceneBooth (Chai et al., 2025) is the first model to integrate both controls in training, combining gated self-attention layers and inpainting ControlNet to train a model based on SD 1.5 (Rombach et al., 2022b) that can accept layout and foreground controls. However, directly combining the two to train a larger model like SDXL can lead to non-convergence or a decline in the quality of generated images. Therefore, we use several strategies when training the layout-controlled inpainting model to ensure training stability, generation quality, and efficiency.
@@ 3. PITA Dataset and PIL Dataset
For automatic product-centric advertising image design, we collect a large-scale Product-Centric Image-Text Advertising (PITA) dataset, gathering 38,017 samples from e-commerce platforms and the CGL dataset (Zhou et al., 2022). 1,000 of them are used for testing. PITA is more diverse than CGL, featuring images with various aspect ratios: 0.684 (51%), 1.0 (33%), 0.667 (14%), and 0.75 (2%). Each image is labeled with prompts (foreground caption and background caption), a product mask, a layout of graphic and nongraphic elements. Each element is represented with a type and a bounding box (bbox). Graphic elements contain “Logo”, “Text”, and “Underlay”. Figure 3 shows some examples in the PITA dataset.
Figure 3. Examples from the PITA dataset, with text, underlay, and logo highlighted in orange, yellow, and red rectangles, respectively.
For layout-controlled background inpainting, we additionally construct a Product-Centric Image Layout (PIL) dataset. The images do not contain graphic elements and labeled with prompts, product masks, and layouts of nongraphic elements. Figure 4 gives some examples. The image scenes cover indoors and outdoors. Product types include furniture, food, clothes, shoes, and so on. PIL contains 12,247 samples in total and 1,000 of them are divided for test.
Figure 4. Examples in PIL dataset.
@@ 4.2. Prompt generation
Training scheme. A standard VLM training scheme involves two stages: pre-training and supervised fine-tuning (SFT) (Liu et al., 2023a). During pre-training, the projector learns to align features between visual and text modalities. In the SFT stage, the projector and LLM are further trained for visual understanding tasks. Since we utilize the VLM trained after the SFT stage, the model can understand input images and follow instructions. We only fine-tune the LLM branch with our dataset to adapt the model to the prompt generation task, using cross-entropy loss as the objective function. The model $\pi_{prompt}$ takes the foreground image $I_{fore}$ and predefined instructions as input, performing two main tasks: generating a foreground description $p_{fore}$ and predicting the background description $p_{back}$. The prompt generation process is as follows:
(1) $p=\pi_{prompt}(I_{fore})$
@@ 4.3. Layout generation
In this stage, we introduce the Jointly Predict Graphic and Nongraphic Layout (JPGNL) method, which aims to optimize the arrangement and improve the readability and appeal of advertising images. It predicts the layout of graphic elements, the product, and other nongraphic elements based on the foreground image, prompt, provided taglines, and target size. Similar to prompt generation, we use XCP2 (Dong et al., 2024) for layout generation. The method is shown in the lower part of Figure 5.
@@
Input data format. For the input to the LLM, we describe the layout generation task and organize input conditions in JSON format. The image input remains the same as in prompt generation. A target canvas aspect ratio is given as condition to support multi-scale layout design. The aspect ratio of the product foreground is provided to help keep its shape. Foreground and background prompts are also given for generating the image layout. For the pre-defined taglines, we list them with empty locations and expect the model to complete them. If a logo is needed, we include it as a condition in the graphic layout section and specify its aspect ratio.
Except for these basic conditions, we also take product characteristics into account. In E-commerce scenarios, some products should not be occluded for better display. While for others, like clothing or human-hold items, partial occlusion is acceptable. To differentiate between these, we introduce a Class-Conditioned Layout Prediction (CCLP) strategy. CCLP provides clear information about the product class and whether foreground occlusion is allowed. For the not-allowing occlusion set (No Occ Set), we describe “The class of subject is [V1]. The bounding boxes of texts should never occlude the subject.” For the allowing occlusion set (Allow Occ Set), the description is “The class of subject is [V2]. The bounding boxes of texts are allowed to occlude the subject.”
Output data format. The output layout is structured in JSON format with three parts: subject layout, other non-graphic layout, and graphic layout. To maintain the shape of the product foreground, we propose a Ratio-Keeping Bbox Representation (RKBR). It is expressed as [x, y, h*r1/r2, h]. (x,y) is the normalized center point. w and h are the normalized width and height. r1 and r2 are the actual aspect ratio of the foreground and canvas respectively. Each value ranges from 0 to 1 with three decimal places. The nongraphic layout contains the location of elements in the background prompt, each having a type (instance name) and bounding box (bbox). The graphic layout defines each element by type, content, and bounding box, covering text, underlay, and logo. The content key is used only for text types to specify its semantic content. Bounding boxes are listed as [x, y, w, h]. The model is expected to supplement underlay adaptively. If logos are required, it is represented in the same way as the product for ratio keeping.
Training scheme. Same as the prompt generation module, we only fine-tune the LLM branch with the visual encoder and projector frozen. The output JSON is structured to predict the layout in a raster scan order, adhering to the ascending order of top-left coordinates. Furthermore, we shuffle the taglines in the input conditional JSON for data augmentation and preventing the model from memorizing the input sequence.

### v2

@@ Abstract.
Creating visually appealing advertising images is often a labor-intensive and time-consuming process. Is it possible to automatically generate such images using only basic product information—specifically, a product foreground image, taglines, and a target size? Existing methods mainly focus on parts of the problem and fail to provide a comprehensive solution. To address this gap, we propose a novel multistage framework called Product-Centric Advertising Image Design (PAID). It consists of four sequential stages to highlight product foregrounds and taglines while achieving overall image aesthetics: prompt generation, layout generation, background image generation, and graphics rendering. Different expert models are designed and trained for the first three stages: First, we use a visual language model (VLM) to generate background prompts that match the products. Next, a VLM-based layout generation model arranges the placement of product foregrounds, graphic elements (taglines and decorative underlays), and various nongraphic elements (objects from the background prompt). Following this, we train an SDXL-based image generation model that can simultaneously accept prompts, layouts, and foreground controls. To support the PAID framework, we create corresponding datasets with over 50,000 labeled images. Extensive experimental results and online A/B tests demonstrate that PAID can produce more visually appealing advertising images.
Figure 1. Generated advertising images by our methods with a product foreground image, taglines, and a target size as input.
@@ 1. Introduction
Advertising images (as shown in Figure 1) are essential for commercial recommendation. Visually appealing and attractive images usually have high click-through rates (Wang et al., 2021) but are labor-intensive and time-consuming to create (Vashishtha et al., 2024). With the rapid advancement of generative methods (Menick and Kalchbrenner, 2018; Van Den Oord et al., 2016; Ho et al., 2020; Rombach et al., 2021), it is now possible to automatically generate advertising images using only basic product inputs: a product foreground image, marketing taglines, and a target size.
Currently, generating advertising images using only the basic inputs mentioned above has rarely been explored. Existing methods (Du et al., 2024; Weng et al., 2024; Wang et al., 2024b; Lin et al., 2023; Zhou et al., 2022) mainly focus on parts of the problem, such as background or layout generation, rather than offering a complete solution for integrating product foregrounds and taglines. Although these methods can be combined, they struggle to effectively highlight product images and taglines while maintaining visual appeal and harmony. Specifically, building on existing approaches (Du et al., 2024; Weng et al., 2024; Wang et al., 2024b; Lin et al., 2023; Zhou et al., 2022), we can first utilize predefined rules to create background prompts. These prompts are then fed into text-to-image inpainting models to complete backgrounds for product foregrounds. Layouts for taglines and decorative elements (collectively referred to as graphic elements) are predicted based on the images and taglines. Finally, attribute prediction and graphics rendering techniques are employed to add these graphic elements to the images. However, this process has several limitations: a) Lack of adaptability: It cannot generate background prompts that match the specific content and shape of the product, leading to poor foreground-background harmony. b) Fixed foreground positioning: The position of the foreground cannot be adjusted based on its characteristics and the background prompt, potentially resulting in inappropriate product size and poor composition. c) Restricted tagline placement: When adding taglines, the whole image content is already set, limiting the available space for tagline placement.
@@
In this paper, we propose a novel framework called Product-Centric Advertising Image Design (PAID) to enhance advertising image generation using only basic product inputs. While PAID remains a multistage process, we have refined its stages and task definitions to enhance foreground-background compatibility and the spatial arrangement of product foregrounds and taglines.
As illustrated in Figure 2, we first introduce a task to generate prompts, using a Visual Language Model (VLM) (Dong et al., 2024) to adaptively create background prompts that better match the scene, placement, and angle based on product foregrounds. Following this, we propose a layout generation task that arranges the graphic elements (marketing taglines and decorative underlays), product foreground, and other nongraphic elements (objects to be generated from the background prompt). This arrangement is based on the product foreground, background prompt, marketing taglines, and target size. By arranging the product foreground and marketing information concurrently, as opposed to generating the overall image first and then deciding the graphic layout, it becomes less constrained by the image content and allows for a more rational presentation of the product info. Furthermore, considering the nongraphic elements together helps minimize conflicts with taglines that could affect readability and attractiveness. Next, we train an inpainting text-to-image model based on SDXL (Podell et al., 2023) that can be controlled by input layouts. Combining inpainting, layout, and prompt control effectively is challenging due to potential conflicts among multiple control signals. A LoRA (Hu et al., 2021) adaption strategy is adopted to balance conflicts and facilitate training. Since layout mainly influences higher-level abstract information, we apply layout control only in the deeper blocks of the UNet (He et al., 2016), combined with a multi-scale training strategy to enhance training stability and efficiency. Finally, we utilize an attribute prediction and graphics rendering module to overlay graphic elements onto the image. To build the framework and verify its effectiveness, we have collected and labeled over 50,000 images to establish datasets. In testing, PAID demonstrates its ability to produce advertising images with better visual quality. It has been deployed on two advertising recommendation scenarios of Taobao, and online A/B tests show that its generated images are more appealing to users.
@@
  • We propose a novel framework PAID for advertising image design given only a product foreground image, taglines, and a target size. It centers on the product foreground and taglines to enhance the overall product prominence and improve the image aesthetics.
@@
  • To arrange product foregrounds and taglines more effectively, a VLM-based layout generator is proposed. It can mix and arrange graphic and nongraphic elements with the target size and prompt. Accordingly, an image generation model that simultaneously accepts layout, prompt, and foreground control is proposed. Several strategies are introduced to enhance effectiveness and efficiency.
  • We have collected and labeled over 50,000 images to create datasets for advertising image generation. The effectiveness of our method has been validated through tests on this dataset and online A/B tests in real advertising scenarios.
@@ 2.1. Advertising Image Design
In recent years, many methods have been proposed to automate advertising image design (Zhou et al., 2022; Hsu et al., 2023; Li et al., 2023b; Seol et al., 2024; Lin et al., 2023; Li et al., 2023a). They can be divided into two subcategories: background-based and foreground-based methods. Background-based methods predict graphic layouts based on complete images (not only product foregrounds), followed by attribute prediction and graphics rendering modules to create the final advertising image. CGL-GAN (Zhou et al., 2022) and DS-GAN (Hsu et al., 2023) design transformer architectures to leverage image info and generate layouts. RADM (Li et al., 2023b) is the first to simultaneously consider the image and tagline content to derive layouts with a diffusion model. PosterLlama (Seol et al., 2024) leverages the rich design knowledge in Large Language Models (LLMs) to predict the layout. AutoPoster (Lin et al., 2023) introduces a complete pipeline which first generates a layout based on a complete image, and then predicts the attributes of graphic elements and renders them onto the image. For these methods, the spatial layout is restricted due to fixed image content, which further restricts the tagline content, including the number and length of taglines.
In contrast, P&R (Li et al., 2023a) is a foreground-based method that leverages both the product foreground image and tagline content to predict the layout of the product foreground and taglines. It subsequently designs a module network to outpaint the background and render the taglines according to the given background prompt and generated layout. The tagline content is not limited, and the layout is more flexible. However, P&R ignores the harmony between the background prompt and the generated layout. The layout may conflict with background generation. In our PAID framework, the overall layout of graphic elements, product foreground, and other nongraphic elements is generated jointly according to the product foreground, background prompt, and tagline contents, aiming to achieve the best harmony among them.
@@ 2.2. Controllable Image Generation
There have been significant developments in text-to-image models (Rombach et al., 2021; Podell et al., 2023). To control instances and further improve image quality, some methods introduce layout guidance, dividing into training-free and training-based categories. Training-free methods manipulate the attention map (Kim et al., 2023) or compute attention-based loss during de-noising (Chefer et al., 2023; Xie et al., 2023; Chen et al., 2024; Bar-Tal et al., 2023). Training-based methods, like GLIGEN (Li et al., 2023c) and InstanceDiffusion (Wang et al., 2024a), modify network architecture, incorporating layout data by adding gated self-attention layers.
For advertising image generation (Du et al., 2024; Wang et al., 2023; Cao et al., 2024; Li et al., 2023a), text-to-image models equipped with inpainting capabilities are frequently utilized. The inpainting technique is employed to preserve product characteristics, while the prompt is used to describe the background. The foreground image is utilized by incorporating extra channels into the UNet architecture (Rombach et al., 2021) or employing an inpainting ControlNet (Zhang et al., 2023). To better control the overall layout of advertising images and improve the quality of generated images, we combine inpainting with layout control. Currently, there is limited exploration in this area. SceneBooth (Chai et al., 2025) is the first model to integrate both controls in training, combining gated self-attention layers and inpainting ControlNet to train a model based on SD 1.5 (Rombach et al., 2021) that can accept layout and foreground controls. However, directly combining the two to train a larger model like SDXL can lead to non-convergence or a decline in the quality of generated images. Therefore, we use several strategies when training the layout-controlled inpainting model to ensure training stability, generation quality, and efficiency.
@@ 3. PITA Dataset and PIL Dataset
For automatic image design of product-centric advertising, we collect a large-scale Product-Centric Image-Tagline Advertising (PITA) dataset, gathering 38,017 samples from e-commerce platforms and the CGL dataset (Zhou et al., 2022). 1,000 of them are used for testing. PITA is a dataset of varying sizes, featuring images with four distinct aspect ratios: 0.684, 1.0, 0.667, and 0.75. Each image is labeled with prompts (foreground caption and background caption), a product mask, and a layout of graphic and nongraphic elements. Each element is represented with a type and a bounding box (bbox). The graphic elements contain “Logo”, “Tagline”, and “underlay”. Some examples of data visualization are in the appendix.
For layout-controlled background inpainting, we additionally construct a Product-Centric Image Layout (PIL) dataset. The images do not contain graphic elements and are labeled with prompts, product masks, and layouts of nongraphic elements. PIL contains 12,247 samples in total and 1,000 of them are divided for test.
@@ 4.2. Prompt generation
Training scheme. A standard VLM training scheme involves two stages: pre-training and supervised fine-tuning (SFT) (Liu et al., 2024a). During pre-training, the projector learns to align features between visual and text modalities. In the SFT stage, the projector and LLM are further trained for visual understanding tasks. Since we utilize the VLM trained after the SFT stage, the model can understand input images and follow instructions. We only fine-tune the LLM branch with our dataset to adapt the model to the prompt generation task, using cross-entropy loss as the objective function. The model $\pi_{prompt}$ takes the foreground image $I_{fore}$ and predefined instructions as input, performing two main tasks: generating a foreground description $p_{fore}$ and predicting the background description $p_{back}$. The prompt generation process is as follows:
(1) $p=\pi_{prompt}(I_{fore}).$
@@ 4.3. Layout generation
In this stage, we introduce the Jointly Predict Graphic and Nongraphic Layout (JPGNL) method, which optimizes the arrangement and improves the tagline readability and image appeal. It predicts the layout of graphic elements, the product, and other nongraphic elements based on the foreground image, prompt, taglines, and target size. Similar to prompt generation, we use XCP2 (Dong et al., 2024) for layout generation. The method is shown in the lower part of Figure 3.
@@
Input data format. For the input to the LLM, we describe the layout generation task and organize input conditions in JSON format. The image input remains the same as in prompt generation. A target canvas aspect ratio is given as a condition to support multi-scale layout design. The aspect ratio of the product foreground is provided to help keep its shape. Foreground and background prompts are also given for generating the image layout. For the pre-defined taglines, we list them with empty locations and expect the model to complete them. If a logo is needed, we include it as a condition in the graphic layout section and specify its aspect ratio.
Except for these basic conditions, we also take product characteristics into account. In E-commerce scenarios, some products should not be occluded for better display. While for others, like clothing or human-hold items, partial occlusion is acceptable. To differentiate between these, we introduce a Class-Conditioned Layout Prediction (CCLP) strategy. CCLP provides clear information about the product class and whether foreground occlusion is allowed. For the not-allowing occlusion set (No Occ Set), we describe “The class of subject is [V1]. The bounding boxes of taglines should never occlude the subject”. For the allowing occlusion set (Allow Occ Set), the description is “The class of subject is [V2]. The bounding boxes of taglines are allowed to occlude the subject”.
Figure 4. Layout control in the image generation model.
Output data format. The output layout is structured in JSON format with three parts: subject layout, other nongraphic layout, and graphic layout. To maintain the shape of the product foreground, we propose a Ratio-Keeping Bbox Representation (RKBR). It is expressed as [x, y, h*r1/r2, h]. (x,y) is the normalized center point. w and h are the normalized width and height. r1 and r2 are the actual aspect ratio of the foreground and canvas, respectively. Each value ranges from 0 to 1 with three decimal places. The nongraphic layout contains the location of elements in the background prompt, each having a type (instance name) and bounding box (bbox). The graphic layout defines each element by type, content, and bounding box, covering tagline, underlay, and logo. The content key is used only for tagline types to specify its semantic content. Bounding boxes are listed as [x, y, w, h]. The model is expected to supplement underlay adaptively. If logos are required, it is represented in the same way as the product for ratio keeping.
Training scheme. Same as the prompt generation module, we only fine-tune the LLM branch with the visual encoder and projector frozen. The output JSON is structured to predict the layout in a raster scan order, adhering to the ascending order of top-left coordinates. Furthermore, we shuffle the taglines in the input conditional JSON for data augmentation and prevent the model from memorizing the input sequence.